# Multimodal Fusion on Low-quality Data: A Comprehensive Survey

### v2

@@ 1 Introduction
Our perception of the world is based on multiple modalities, e.g., touch, sight, hearing, smell and taste. Even when some sensory signals are unreliable, humans can extract useful clues from imperfect multimodal inputs and further piece together the whole scene of happening events [1]. With the development of sensory technology, we can easily collect diverse forms of data for analysis. To fully release the value of each modality, multimodal fusion emerges as a promising paradigm to obtain precise and reliable predictions by integrating all available cues for downstream analysis tasks, e.g., medical image analysis, autonomous vehicles [2, 3] and sentiment recognition [4, 5, 6]. Intuitively, fusing information from different modalities offers the possibility of exploring cross-modal correlation and gaining better performance. However, there is growing recognition that widely-used AI models are often misled by spurious correlations and biases within low-quality data. In real-world setting, the quality of different modalities usually varies due to unexpected environmental factors or sensor issues. Some recent studies have shown both empirically and theoretically that conventional multimodal fusion may fail on low-quality multimodal data in the wild, e.g., imbalanced [7, 8, 9, 10], noisy [11] or even corrupted [12] multimodal data. To overcome such a limitation and take a step to strong and generalized multimodal learning in the real world, we identify the properties of low-quality multimodal data and focus on some unique challenges for multimodal machine fusion in real-world setting. We also highlight technical advances that may help to make multimodal fusion more reliable and trustworthy in the open environment. In this paper, we identify and explore four core technical challenges surrounding multimodal fusion on low-quality multimodal data. They are summarized as follows (also intuitively illustrated in Fig. 1):
@@
Figure 1: Illustrations of challenges for machine learning on low-quality multimodal data. Blue and gold represent various modalities. Deeper color denotes higher quality.
TABLE I: A summary of the notations we mainly used in multimodal fusion .
$\mathbf{x}=[x^{1},\cdots,x^{m}]$ multimodal input consists of $m$ modalities
$\mathcal{X}=\mathcal{X}^{1}\times\cdots\times\mathcal{X}^{m}$ multimodal input space consists of $m$ unimodal space
$y\in\mathcal{Y}$ target $y$ and corresponding space $\mathcal{Y}$
$\mathbf{z}\in\mathcal{Z}$ multimodal latent representation $\mathbf{z}$ and latent space $\mathcal{Z}$
$h:\mathcal{X}\rightarrow\mathcal{Z}$ multimodal fusion mapping
$g:\mathcal{Z}\rightarrow\mathcal{Y}$ task-specific mapping
@@ 2 Learning on Noisy Multimodal Data
Collecting high-quality multimodal data in real-world scenarios inevitably faces significant challenges due to the presence of noise. Multimodal data [15] noise can arise from sensor errors [16], environmental interference, or transmission losses. For vision modality, the electronic noise in sensors leads to detail loss. Besides, audio modality may suffer from unexpected distortion due to environmental factors. To make things worse, weakly aligned or even unaligned multimodal samples are also commonly presented, which lies in a higher-level semantic space. Fortunately, considering the correlation between multiple modalities or better utilizing multimodal data can help the fusion of noisy multimodal data. Various related works [17, 18, 16] demonstrate that multimodal models surpass their unimodal counterparts. This can be attributed to the ability of multimodal data to identify and mitigate potential noise by leveraging the correlations among different modalities.
@@ 2.1 Modal-specific noise reduction
2.1.1 Weighted average fusion
A simple multimodal noise reduction way is to conduct average fusion over multimodal data. Due to the randomness of noise, the average operation effectively reduces the ratio of noise in fused images. Rajalingam et al. [19] developed an intelligent multimodal fusion algorithm. The method decomposes the input image into high-frequency and low-frequency components and proposes an average fusion rule for the low-frequency components while utilizing guided filtering for the high-frequency components. Essentially, average-based multimodal noise reduction methods assign the same fusion weights to different modalities. However, the noise severity of each modality is different and varies among multimodal samples. To address the aforementioned issue, weighted average-based multimodal noise reduction method is proposed. Budhiraja et al. [20] proposed a multimodal image fusion scheme based on the guided filter. Their approach involves Gaussian decomposition with local average energy and average gradient-based salient maps for the base layer and detail layer, respectively. Qu et al. [21] proposed a fusion rule used for calculating the wavelet transformation modulus maxima of input images with different bandwidths and levels. The approach aims to effectively combine the information from multiple image modalities. Achim et al.[22] developed a fusion rule based on fractional lower-order moments for magnetic resonance (MR) and Computed Tomography (CT) images in the multi-scale medical image analysis literature. Considering the similarities and disparities between the different modalities, Yu et al[23] divided the image into the common component and innovation component and represented the source image with the common and innovation sparse coefficients by joint sparse representation. Wang et al. [24] designed a gate mechanism to dynamically fuse visual features. By utilizing visual features as supplementary information, the proposed method effectively mitigates the influence of noise in voice signals. However, this strategy assumes that the visual data is clean. If the visual data is also corrupted by noise, this solution may be sub-optimal. Sadeghi et al. [25] adopted a latent sequential variable with Markovian dependencies to decide whether visual data is beneficial to audio-visual speech enhancement. Bai et al.[18] proposed a soft-association mechanism to handle inferior image conditions. In practice, the self-attention in Transformer can be used to adaptively decide where and what information should be taken from the image. Cheng et al. [16] adopted both soft attention and hard attention to mitigate noise issue, which is caused by rolling-shutter effect in lidar-stereo fusion task. Specifically, a fusion block is utilized to aggregate supplement information from stereo cameras and generates two disparity maps. Then, by comparing disparity map and lidar point, the noisy points in raw data are filtered out.
@@ 2.2 Cross-modal noise reduction
Many multimodal tasks (e.g., multimodal object detection, vision-and-language understanding) highly depends on correctly aligned multimodal training data. However, the real-world multimodal pairs tend to contain weakly aligned or even unaligned samples [32]. For example, in RGB/thermal multimodal object detection, multimodal inputs are often weakly aligned, i.e., the locations of the same objects may be shifted across different modalities [33]. In social media, the text descriptions are frequently irrelevant with the image content, i.e., unpaired modalities. In this section, we consider weakly aligned or unaligned multimodal samples as cross-modal noise. Compared to modal-specific noise, cross-modal noise lies in a higher-level semantic space. Current cross-modal noise reduction methods can be roughly split into rules-based filtering, model-based rectifying and noise robustness regularization.
From the data perspective, some strict rules are used to conduct data clearning [34, 35, 36]. As a recent representative method, CAT (Complexity, Action, and Text-spotting) is a filtering strategy which is devised to select informative image-text pairs, thereby reducing the influence of cross-modal noise [35]. In multispectral object detection, image registration (i.e., spatial alignment) is a commonly used preprocessing [37, 38]. By employing geometric rules to align two images, the shifted positions can be rectified across different modalities.
@@ 2.3 Discussion
TABLE II: A summary of representative multi-modal noise reduction works.
Noise type Denoise Paradigm Application Reference
Average fusion Image fusion [19]
Modal-specific Weighted fusion [22]
[20]
Audio enhancement [24] [25]
Object detection [18]
Lidar-stereo fusion [16]
RGB-D object recognition [44]
Image classification [45]
Representation learning [46]
Joint variation Image fusion [27]
[47]
[28]
[29]
[31]
Remote-sensing scene classification [30]
Cross-modal Rules-based filtering Multimodal pretraining [34, 35, 36]
Model-based rectifying Image-text matching [12]
Multimodal pretraining [39, 40]
Noise robust regularization Multimodal pretraining [41, 43]
Vision-language understanding [42]
@@ 3 Incomplete multimodal Learning
The collected multimodal data in real applications are often incomplete with partial missing modalities for some samples due to unexpected factors, such as equipment damage, data transmission and storage loss. For a concrete example, in user-oriented recommendation system, the browsing behavior history and credit score information may not always be available for some users [48]. Similarly, although combining data from multiple modalities e.g., magnetic resonance imaging (MRI) scans, positron emission tomography (PET) scans, and cerebrospinal fluid (CSF) information, can obtain a more accurate diagnosis for Alzheimer’s Disease [49, 50], owing to high measurement cost for PET scans and uncomfortable invasive tests for CSF, some patients may refuse to conduct these examinations. Thus, incomplete multimodal data commonly exist in Alzheimer’s Disease diagnosis [51]. Typically, conventional multimodal learning models assume the completeness of multimodal data and thus cannot be applied to partial missing modalities directly. For this issue, incomplete multimodal learning, aiming at exploring information of incomplete multimodal data with partial missing modalities, appears and has received increasing research interests in recent years [52]. In this section, we mainly focus on the current progress of researches on incomplete multimodal learning. From the view of whether to impute the missing data, we categorize existing methods into two groups including imputation-based and imputation-free incomplete multimodal learning, where the imputation based methods are further divided into two groups as shown in Fig. 2 including instance and modality level imputation.
@@ 3.1.2 Learning based imputation
Kernel/graph-based imputation: Kernel and graph learning based approaches are two main lines of incomplete multimodal learning, which transform the learning phase from the raw data to the kernel space or graph. For the kernel and graph constructed from each modality, each element represents the relationship in kernel space or affinity/similarity of instance pairs. The partial missing instances in each modality result in incomplete kernels and graphs. A line of works has been proposed to address the incomplete kernel/graph-based multimodal learning issue. Here we highlight a few seminal works. David Williams et al[59] proposed a Gaussian mixture model based method that can analytically compute the missing Gaussian kernel elements associated with the missing instances. Trivedi et al[60] proposed a kernelized canonical correlation analysis (KCCA) based imputation method, which minimizes a Laplacian kernel regularization to obtain the missing kernel elements. Although effective, it is only applicable to the two-modal data with one complete modality. Lately, collective kernel learning (CKL) [48] is proposed to mutually complete the two incomplete kernels associated with the two incomplete modalities. For the data with more than two modalities, such as the case with $k$ modalities ($k\geq 3$), CKL needs to recover each kernel matrix ${K_{i+1}}$ of the $\left({i+1}\right)$th modality according to kernel ${K_{i}}$ in a cycle iteration scheme. Sahely et al[61] proposed a multimodal kernel completion method based on the sparse reconstruction technique. The method can complete the missing rows and columns of all kernels jointly in a within-kernel preservation and between-kernel exploration framework. Yang et al[62] reduced incomplete multi-modal imputation into semi-supervised learning problem. They further proposed SLIM (Semi-supervised Learning with Incomplete Modality ) and its kernel vision SLIM-K based on matrix completion for solving the modal incompleteness. Recently, Liu et al. proposed several flexible kernel methods to recover the missing kernel elements for incomplete multimodal clustering tasks [63, 64, 65, 66]. For instance, MKKM-IK [63] unifies the kernel completion and consensus partition matrix learning in a unified framework:
$\begin{array}[]{l}\mathop{\min}\limits_{H,\gamma,\left\{{{K^{\left(m\right)}}}\right\}_{m=1}^{l}}\text{\rm Tr}\left({\sum\limits{\gamma_{m}^{2}{K^{\left(m\right)}}\left({I-H{H^{T}}}\right)}}\right)\\ s.t.HH=I,\sum\nolimits_{m=1}^{l}{{\gamma_{m}}}=1,0\leq\gamma\leq 1,\\ \forall\left\{{i,j}\right\}\in{\Omega^{\left(m\right)}},K_{i,j}^{\left(m\right)}={\kappa^{(m)}}\left({x_{i}^{\left(m\right)},x_{j}^{\left(m\right)}}\right),\end{array}$ (3)
where ${\kappa^{\left(m\right)}}$ denotes the kernel function for the $m$-th modality. ${\Omega^{\left(m\right)}}$ denotes the set of location index corresponding to the un-missing instance in the $m$-th modal. ${K^{\left(m\right)}}\in{R^{n\times n}}$ represents the kernel of the $m$-th modality to recover, in which some elements $K_{i,j}^{\left(m\right)}$ corresponding to the observed instances in the $m$-th modality are unchanged during the optimization. $H\in{R^{n\times c}}$ is the consensus partition matrix for data clustering.
@@
To obtain a more reliable modality imputation performance, Tang et al[74] proposed to impute the modality as the average of corresponding semantic neighbors. Lin et al[75] established a dual prediction model that can predict the latent representation of the missing modality according to that of the observed modality and then used the decoder to obtain the recovered missing modality according to the predicted latent representation. matrix factorization is also applied for raw modality imputation. For example, UEAF (Unified Embedding Alignment Framework) [76] and IMCRV (Incomplete multimodal Clustering with Reconstructed Views) [77] provide two matrix factorization based missing modality recovery models, which can reversely reconstruct the data of missing modalities from the common representation. In [78], low-rank representation and tensor learning are introduced for joint missing modality recovery and graph completion. Explicitly taking the reliability of imputation into account, Xie et al[79] introduced the Uncertainty-induced Incomplete Multi-View Data Classification (UIMC) model to address challenges in classifying incomplete multi-view data with missing perspectives. UIMC explores and exploits uncertainty by imputing missing modalities as distributions rather than deterministic values, allowing for more reliable predictions. The model employs an evidential multimodal learning framework to weigh imputed data, reducing the impact of low-quality imputations.
@@ 3.2.1 Latent representation and projection learning
Latent representation and projection learning based incomplete multimodal learning (taking clustering for example) generally seek to obtain a multimodal latent representation by exploring the available partially aligned information among these available modalities. In this branch, partial multimodal clustering (PMVC) [85] is a popular and pioneering work based on matrix factorization. It concentrates on encoding the data with two fully observed modalities into a common representation within the latent subspace. Based on PMVC, incomplete multimodal grouping (IMG) [54] and partial multimodal subspace clustering [86] further introduce the graph constraints to capture the structure information in the latent common subspace. A basic model used in the above methods to learn the latent common representation from the incomplete bi-modal data can be formulated as follows:
$\begin{array}[]{l}\mathop{\min}\limits_{{P_{o}},\left\{{P_{uo}^{\left(m\right)},{U^{\left(m\right)}}}\right\}_{m=1}^{2}}\sum\limits_{m=1}^{2}{\left\|{\left[\begin{array}[]{l}X_{o}^{\left(m\right)}\\ X_{uo}^{\left(m\right)}\end{array}\right]-\left[\begin{array}[]{l}{P_{o}}\\ P_{uo}^{\left(m\right)}\end{array}\right]{U^{\left(m\right)}}}\right\|_{F}^{2}}\\ {\kern 80.0pt}+\Psi\left({{P_{o}},P_{uo}^{1},P_{uo}^{2},{U^{1}},{U^{2}}}\right),\end{array}$ (4)
where ${X^{(1)}}=\left[\begin{array}[]{l}X_{o}^{(1)}\\ X_{uo}^{(1)}\end{array}\right]$ and ${X^{(2)}}=\left[\begin{array}[]{l}X_{o}^{(2)}\\ X_{uo}^{(2)}\end{array}\right]$ denote the data in the two modalities, respectively. ${X_{o}}=\left\{{X_{o}^{(1)},X_{o}^{(2)}}\right\}$ denotes the samples with two fully observed modalities. $X_{uo}^{(1)}$ and $X_{uo}^{(2)}$ are the data with only the first modality and second modality, respectively. $U^{1}$ and $U^{(2)}$ are the basis matrices. $\Psi\left({{P_{o}},P_{uo}^{(1)},P_{uo}^{(2)},{U^{(1)}},{U^{(2)}}}\right)$ denotes the constraints on these variables. The learned common representation can be represented as $P=\left[{{P_{o}};P_{uo}^{(1)};P_{uo}^{(2)}}\right]$.
Zhou et al[87] integrated the latent representation learning and regression into a unified framework for Alzheimer’s disease diagnosis based on incomplete multimodal neuroimaging and genetic data. The motivation of the latent representation learning module for incomplete multimodal data is similar to model (4). However, a widely recognized limitation of the above methods based on model (4) is that they are only applicable to the case with bi-modal data. To address this limitation, weighted matrix factorization is introduced, where doubly aligned incomplete multimodal clustering [88], one-pass incomplete multimodal clustering [89], and localized sparse incomplete multimodal clustering [90] are representative methods. These models generally impose the location information of missing instances as the weighted matrices with binary values on the matrix factorization model to eliminate the loss yet reduce the negative influence of the missing instances. The basic model of the weighted matrix factorization based methods can be formulated as follows:
@@
Compared with the basic model (4), the methods based on model (5) are more flexible and can be easily extended to the data with more than two modalities and arbitrary incomplete cases. Besides of these weighted matrix factorization models, a generalized incomplete multimodal clustering model is proposed to learn a complete common representation from the incomplete modal-specific representations with a partially aligned learning term [91]. Specifically, such a partially aligned learning term can be regarded as a variant of the weighted common representation learning. To prevent the model from overly focusing on certain modalities with larger dimensions, projective incomplete multimodal clustering transforms the matrix factorization into a variant of the weighted projection learning model to obtain the consensus representation from the incomplete multimodal data [92]. For the incomplete multimodal and incomplete multi-label classification task, a variant of the weighted projection learning model is developed to explore the information of available modalities [93]. In terms of infant neuroscience, Zhang et al. propose to explore incomplete multimodal data by constructing a latent representation to capture the complementary information from incomplete time-points [94]. In summary, the above methods focus on exploring the available information of the observed modalities by introducing some techniques like the weighting mechanism to eliminate the loss of those missing modalities.

### v3

@@ 1 Introduction
Our perception of the world is based on multiple modalities, e.g., touch, sight, hearing, smell and taste. Even when some sensory signals are unreliable, humans can extract useful clues from imperfect multimodal inputs and further piece together the whole scene of happening events [1]. With the development of sensory technology, we can easily collect diverse forms of data for analysis. To fully release the value of each modality, multimodal fusion emerges as a promising paradigm to obtain precise and reliable predictions by integrating all available cues for downstream analysis tasks, e.g., medical image analysis, autonomous vehicles [2, 3] and sentiment recognition [4, 5, 6]. Intuitively, fusing information from different modalities offers the possibility of exploring cross-modal correlation and gaining better performance. However, there is growing recognition that widely-used AI models are often misled by spurious correlations and biases within low-quality data. In real-world setting, the quality of different modalities usually varies due to unexpected environmental factors or sensor issues. Some recent studies have shown both empirically and theoretically that conventional multimodal fusion may fail on low-quality multimodal data in the wild, e.g., imbalanced [7, 8, 9, 10], noisy [11] or even corrupted [12] multimodal data. To overcome such a limitation and take a step to strong and generalized multimodal learning in the real world, we identify the properties of low-quality multimodal data and focus on some unique challenges for multimodal machine fusion in real-world setting. We also highlight technical advances that may help to make multimodal fusion more reliable and trustworthy in the open environment. In this paper, we identify and explore four core technical challenges surrounding multimodal fusion on low-quality multimodal data. They are summarized as follows:
@@
Figure 1: Illustrations of challenges for machine learning on low-quality multimodal data. Blue and gold represent various modalities. Deeper color denotes higher quality. Assuming we have $N$ multimodal samples consisting $M$ different modalities, the dimension of each modality is $D$. $q(x)$ denotes the quality of multimodal input, i.e., the information collected from $x$ that can support the downstream tasks. (a) The quality of noisy multimodal data is randomly influenced by unexpected environmental factors. (b) Certain modalities of incomplete multimodal data are of zero quality (do not convey any useful information). (c) The expected quality of modalities are different for imbalanced multimodal data. (d) The quality of different modality are varying for samples.
@@ 2 Learning on Noisy Multimodal Data
Collecting high-quality multimodal data in real-world scenarios inevitably faces significant challenges due to the presence of noise. Multimodal data [15] noise can arise from sensor errors [16], environmental interference, or transmission losses. For vision modality, the electronic noise in sensors leads to detail loss. Besides, audio modality may suffer from unexpected distortion due to environmental factors. To make things worse, weakly aligned or even unaligned multimodal samples are also commonly presented, which can be seen as more complex noise that lies in a higher-level semantic space. Fortunately, considering the correlation between multiple modalities or better utilizing multimodal data can help the fusion of noisy multimodal data. Various related works [17, 18, 16] demonstrate that multimodal models surpass their unimodal counterparts. This can be attributed to the ability of multimodal data to identify and mitigate potential noise by leveraging the correlations among different modalities.
@@ 2.2 Cross-modal noise reduction
Many multimodal tasks (e.g., multimodal object detection, vision-and-language understanding) highly depends on correctly aligned multimodal training data. However, the real-world multimodal pairs tend to contain weakly aligned or even unaligned samples [25]. For example, in RGB-thermal multimodal object detection, multimodal inputs are often weakly aligned, i.e., the locations of the same objects may be shifted across different modalities [26]. In social media, the text descriptions are frequently irrelevant with the image content, i.e., unpaired modalities. In this section, we consider weakly aligned or unaligned multimodal samples as cross-modal noise. Compared to modal-specific noise, cross-modal noise lies in a higher-level semantic space. Current cross-modal noise reduction methods can be roughly split into rules-based filtering, model-based rectifying and noise robustness regularization.
From the data perspective, some strict rules are used to conduct data cleaning [27, 28, 29]. As a recent representative method, CAT (Complexity, Action, and Text-spotting) is a filtering strategy which is devised to select informative image-text pairs, thereby reducing the influence of cross-modal noise [28]. In multi-spectral object detection, image registration (i.e., spatial alignment) is a commonly used pre-processing [30, 31]. By employing geometric rules to align two images, the shifted positions can be rectified across different modalities.
@@ 2.3 Discussion
TABLE I: A summary of representative multi-modal noise reduction works.
Noise type Reference Denoise Paradigm Application Modal-specific [20] Joint variation Image fusion [37] [21] [22] [24] [23] Remote-sensing scene classification Cross-modal [27, 28, 29] Rules-based filtering Multimodal pretraining [12] Model-based rectifying Image-text matching [32, 33] Multimodal pretraining [34, 36] Noise robust regularization Multimodal pretraining [35] Vision-language understanding
@@ 3 Incomplete multimodal Learning
The collected multimodal data in real applications are often incomplete with partial missing modalities for some samples due to unexpected factors, such as equipment damage, data transmission and storage loss. For a concrete example, in recommendation system, the browsing behavior history and credit score information may not always be available for some users [38]. Similarly, multimodal data can be highly incomplete in medical analysis due to the patients’ personal preferences, economic constraints and limication in medical resources. For example, although combining data from multiple modalities e.g., magnetic resonance imaging (MRI) scans, positron emission tomography (PET) scans, and cerebrospinal fluid (CSF) information can obtain a more accurate diagnosis for Alzheimer’s Disease [39, 40], owing to high measurement cost for PET scans and uncomfortable invasive tests for CSF, some patients may refuse to conduct these examinations. Thus, incomplete multimodal data commonly exist in Alzheimer’s Disease diagnosis [41]. Typically, conventional multimodal learning models assume the completeness of multimodal data and thus cannot be applied to partial missing modalities directly. For this issue, incomplete multimodal learning, aiming at exploring information of incomplete multimodal data with partial missing modalities, appears and has received increasing research interests in recent years [42]. In this section, we mainly focus on the current progress of researches on incomplete multimodal learning. From the view of whether to impute the missing data or not, we categorize existing methods into two groups including imputation-based and imputation-free incomplete multimodal learning, where the imputation based methods are further divided into two groups as shown in Fig. 2 including instance and modality level imputation.
@@ 3.1.2 Learning based imputation
Kernel/graph-based imputation: Kernel and graph learning based approaches are two main lines of incomplete multimodal learning, which transform the learning phase from the raw data to the kernel space or graph. For the kernel and graph constructed from each modality, each element represents the relationship in kernel space or affinity/similarity of instance pairs. The partial missing instances in each modality result in incomplete kernels and graphs. A line of works has been proposed to address the incomplete kernel/graph-based multimodal learning issue. Here we highlight a few seminal works. David Williams et al[49] proposed a Gaussian mixture model based method that can analytically compute the missing Gaussian kernel elements associated with the missing instances. Trivedi et al[50] proposed a kernelized canonical correlation analysis (KCCA) based imputation method, which minimizes a Laplacian kernel regularization to obtain the missing kernel elements. Although effective, it is only applicable to the two-modal data with one complete modality. Lately, collective kernel learning (CKL) [38] is proposed to mutually complete the two incomplete kernels associated with the two incomplete modalities. For the data with more than two modalities, such as the case with $k$ modalities ($k\geq 3$), CKL needs to recover each kernel matrix ${K_{i+1}}$ of the $\left({i+1}\right)$th modality according to kernel ${K_{i}}$ in a cycle iteration scheme. Sahely et al[51] proposed a multimodal kernel completion method based on the sparse reconstruction technique. The method can complete the missing rows and columns of all kernels jointly in a within-kernel preservation and between-kernel exploration framework. Yang et al[52] reduced incomplete multi-modal imputation into semi-supervised learning problem. They further proposed SLIM (Semi-supervised Learning with Incomplete Modality ) and its kernel vision SLIM-K based on matrix completion for solving the modal incompleteness. Recently, Liu et al. proposed several flexible kernel methods to recover the missing kernel elements for incomplete multimodal clustering tasks [53, 54, 55, 56].
@@
To obtain a more reliable modality imputation performance, Tang et al[64] proposed to impute the modality as the average of corresponding semantic neighbors. Lin et al[65] established a dual prediction model that can predict the latent representation of the missing modality according to that of the observed modality and then used the decoder to obtain the recovered missing modality according to the predicted latent representation. matrix factorization is also applied for raw modality imputation. For example, UEAF (Unified Embedding Alignment Framework) [66] and IMCRV (Incomplete multimodal Clustering with Reconstructed Views) [67] provide two matrix factorization based missing modality recovery models, which can reversely reconstruct the data of missing modalities from the common representation. In [68], low-rank representation and tensor learning are introduced for joint missing modality recovery and graph completion. Explicitly taking the reliability of imputation into account, Xie et al[69] introduced the Uncertainty-induced Incomplete Multi-View Data Classification (UIMC) model to address challenges in classifying incomplete multi-view data with missing perspectives. UIMC explores and exploits uncertainty by imputing missing modalities as distributions rather than deterministic values, allowing for more reliable predictions. The model employs an evidential multimodal learning framework to weigh imputed data, reducing the impact of low-quality imputations. In incomplete multimodal action recognition task, Sangmin et al. proposed a modular network called ActionMAE, which randomly drops modality features and then learn to impute them.
@@ 3.2.1 Latent representation and projection learning
Latent representation and projection learning based incomplete multimodal learning (taking clustering for example) generally seek to obtain a multimodal latent representation by exploring the available partially aligned information among these available modalities. In this branch, partial multimodal clustering (PMVC) [75] is a popular and pioneering work based on matrix factorization. It concentrates on encoding the data with two fully observed modalities into a common representation within the latent subspace. Based on PMVC, incomplete multimodal grouping (IMG) [44] and partial multimodal subspace clustering [76] further introduce the graph constraints to capture the structure information in the latent common subspace. A basic model used in the above methods to learn the latent common representation from the incomplete bi-modal data.
Zhou et al[77] integrated the latent representation learning and regression into a unified framework for Alzheimer’s disease diagnosis based on incomplete multimodal neuroimaging and genetic data. However, a widely recognized limitation of the above methods based on model (4) is that they are only applicable to the case with bi-modal data. To address this limitation, weighted matrix factorization is introduced, where doubly aligned incomplete multimodal clustering [78], one-pass incomplete multimodal clustering [79], and localized sparse incomplete multimodal clustering [80] are representative methods. These models generally impose the location information of missing instances as the weighted matrices with binary values on the matrix factorization model to eliminate the loss yet reduce the negative influence of the missing instances. The basic model of the weighted matrix factorization based methods can be formulated as follows:
@@
Besides of these weighted matrix factorization models, a generalized incomplete multimodal clustering model is proposed to learn a complete common representation from the incomplete modal-specific representations with a partially aligned learning term [81]. Specifically, such a partially aligned learning term can be regarded as a variant of the weighted common representation learning. To prevent the model from overly focusing on certain modalities with larger dimensions, projective incomplete multimodal clustering transforms the matrix factorization into a variant of the weighted projection learning model to obtain the consensus representation from the incomplete multimodal data [82]. For the incomplete multimodal and incomplete multi-label classification task, a variant of the weighted projection learning model [83] and weighted matrix factorization model [84] are developed to explore the information of available modalities. In terms of infant neuroscience, Zhang et al. propose to explore incomplete multimodal data by constructing a latent representation to capture the complementary information from incomplete time-points [85]. In summary, the above methods focus on exploring the available information of the observed modalities by introducing some techniques like the weighting mechanism to eliminate the loss of those missing modalities.